# Leveraging Dual Process Theory in Language Agent Framework for Real-time Simultaneous Human-AI Collaboration

### v3

@@ Abstract
Agents built on large language models (LLMs) have excelled in turn-by-turn human-AI collaboration but struggle with simultaneous tasks requiring real-time interaction. Latency issues and the challenge of inferring variable human strategies hinder their ability to make autonomous decisions without explicit instructions. Through experiments with current independent System 1 and System 2 methods, we validate the necessity of using Dual Process Theory (DPT) in real-time tasks. We propose DPT-Agent, a novel language agent framework that integrates System 1 and System 2 for efficient real-time simultaneous human-AI collaboration. DPT-Agent’s System 1 uses a Finite-state Machine (FSM) and code-as-policy for fast, intuitive, and controllable decision-making. DPT-Agent’s System 2 integrates Theory of Mind (ToM) and asynchronous reflection to infer human intentions and perform reasoning-based autonomous decisions. We demonstrate the effectiveness of DPT-Agent through further experiments with rule-based agents and human collaborators, showing significant improvements over mainstream LLM-based frameworks. DPT-Agent can effectively help LLMs convert correct slow thinking and reasoning into executable actions, thereby improving performance. To the best of our knowledge, DPT-Agent is the first language agent framework that achieves successful real-time simultaneous human-AI collaboration autonomously. Code of DPT-Agent can be found in https://github.com/sjtu-marl/DPT-Agent.
@@
Shao Zhang1thanks: Equal Contribution, Xihuai Wang [^1] footnotemark: 1thanks: Work done while interning at Meituan., Wenhao Zhang1, Chaoran Li1, Junru Song1, Tingyu Li1, Lin Qiu2, Xuezhi Cao2, Xunliang Cai2, Wen Yao3, Weinan Zhang1, Xinbing Wang1, Ying Wen1thanks:  Corresponding Author: ying.wen@sjtu.edu.cn. 1Shanghai Jiao Tong University, 2Meituan, 3Intelligent Game and Decision Laboratory
@@ 4.2.2 Finite-state Machine & Action Executor
When LLM generates code-as-policy, the executable code changes the pre-defined logics of FSM and thus facilitates the adaption to human and performance improvement. The FSM takes in the code-as-policy and game states, and outputs macro actions, denoted as $ma$, which are high-level combinations of atomic actions for specific targets. For example, in Overcooked, macro actions include food ingredients preparation, food assembling and food serving. The generated macro actions are sent to an action executor for conversion into atomic actions that can be directly executed in the environment. The action executor employs script policies, ensuring smooth and efficient execution. Upon receiving a macro action, the action executor selects an appropriate execution plan and performs path planning to determine the necessary atomic actions using the A* algorithm Hart et al. (1968). The Detailed design and implementation of the FSM is provided in Section A.1. These processes can be formalized as $ma_{t}=\text{FSM}(c_{t},s_{t})$ and $a_{t}^{i}=\text{Executor}\left(ma_{t}\right)~{}$.
@@ 5.1 Overcooked Challenge for Real-time Simultaneous Human-AI Collaboration
Figure 5: Results of LLM with ReAct, Refelxion and DPT-Agent w/o ToM in the Single Agent Game.
@@ 6.1 Capability in Real-time Task
ReAct Refelxion DPT-Agent DPT-Agent ReAct Refelxion DPT-Agent DPT-Agent
@@ Appendix D Implementation of Act, ReAct and Reflexion Frameworks
We implement Act, ReAct and Refelxion in Overcooked challenge. The three frameworks use the same prompt in the instruction part with DPT-Agent in Section C.1. We outline the specific differences in the output prompts for the three frameworks below.
@@ G.2 Detailed Results
Table 6: Performance of Different Models in LLM as Indenpendent System 1 and System 2 Experiments.
@@
ReAct Refelxion DPT-Agent DPT-Agent ReAct Refelxion DPT-Agent DPT-Agent
@@ I.1 Procedure
In the formal experiment, after each game, participants were asked to rank the agents based on their collaborative capabilities and personal preference. Upon completing all eight games, they filled out an additional questionnaire, where we collected their perceptions of task load and their intended level of task engagement.

### v5

@@ Abstract
Agents built on large language models (LLMs) have excelled in turn-by-turn human-AI collaboration but struggle with simultaneous tasks requiring real-time interaction. Latency issues and the challenge of inferring variable human strategies hinder their ability to make autonomous decisions without explicit instructions. Through experiments with current independent System 1 and System 2 methods, we validate the necessity of using Dual Process Theory (DPT) in real-time tasks. We propose DPT-Agent, a novel language agent framework that integrates System 1 and System 2 for efficient real-time simultaneous human-AI collaboration. DPT-Agent’s System 1 uses a Finite-state Machine (FSM) and code-as-policy for fast, intuitive, and controllable decision-making. DPT-Agent’s System 2 integrates Theory of Mind (ToM) and asynchronous reflection to infer human intentions and perform reasoning-based autonomous decisions. We demonstrate the effectiveness of DPT-Agent through further experiments [^1] The up-to-date evaluation results will be maintained in AGI-Eval: Overcooked Challenge. with rule-based agents and human collaborators, showing significant improvements over mainstream LLM-based frameworks. DPT-Agent can effectively help LLMs convert correct slow thinking and reasoning into executable actions, thereby improving performance. To the best of our knowledge, DPT-Agent is the first language agent framework that achieves successful real-time simultaneous human-AI collaboration autonomously. Code of DPT-Agent can be found in https://github.com/sjtu-marl/DPT-Agent.
@@
Shao Zhang1thanks: Equal Contribution, Xihuai Wang [^1] footnotemark: 1thanks: Work done while interning at Meituan., Wenhao Zhang1, Chaoran Li1, Junru Song1, Tingyu Li1, Lin Qiu2, Xuezhi Cao2, Xunliang Cai2, Wen Yao3, Weinan Zhang1, Xinbing Wang1, Ying Wen1thanks:  Corresponding Author. 1Shanghai Jiao Tong University, 2Meituan, 3Intelligent Game and Decision Laboratory {shaozhang,leoxhwang,ying.wen}@sjtu.edu.cn
@@ 4.2.2 Finite-state Machine & Action Executor
When LLM generates code-as-policy, the executable code changes the pre-defined logic of FSM and thus facilitates the adaption to human and performance improvement. The FSM takes in the code-as-policy and game states, and outputs macro actions, denoted as $ma$, which are high-level combinations of atomic actions for specific targets. For example, in Overcooked, macro actions include food ingredients preparation, food assembling and food serving. The generated macro actions are sent to an action executor for conversion into atomic actions that can be directly executed in the environment. The action executor employs script policies, ensuring smooth and efficient execution. Upon receiving a macro action, the action executor selects an appropriate execution plan and performs path planning to determine the necessary atomic actions using the A* algorithm Hart et al. (1968). The Detailed design and implementation of the FSM is provided in Section A.1. These processes can be formalized as $ma_{t}=\text{FSM}(c_{t},s_{t})$ and $a_{t}^{i}=\text{Executor}\left(ma_{t}\right)~{}$.
@@ 5.1 Overcooked Challenge for Real-time Simultaneous Human-AI Collaboration
Figure 5: Results of LLM with ReAct, Reflexion and DPT-Agent w/o ToM in the Single Agent Game.
@@ 6.1 Capability in Real-time Task
ReAct Reflexion DPT-Agent DPT-Agent ReAct Reflexion DPT-Agent DPT-Agent
@@ 8 Conclusion
Acknowledgments
Team from Shanghai Jiao Tong University is supported by the National Key R&D Program of China (2024YFC3505402), National Natural Science Foundation of China (T2421002, 62061146002, 62020106005, U2244217, 62322603), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and Shanghai Pilot Program for Basic Research - Shanghai Jiao Tong University. Xihuai Wang is supported by the Wen-Tsun Wu AI Honorary Doctoral Scholarship from AI Institute, Shanghai Jiao Tong University. We extend heartfelt thanks to the participants in our human experiments from Shanghai Jiao Tong University. We would like to express our gratitude to the AGI-Eval community for their support in hosting the evaluation results of this work.
@@ Appendix D Implementation of Act, ReAct and Reflexion Frameworks
We implement Act, ReAct and Reflexion in Overcooked challenge. The three frameworks use the same prompt in the instruction part with DPT-Agent in Section C.1. We outline the specific differences in the output prompts for the three frameworks below.
@@ G.2 Detailed Results
QwQ-32b 0.55(0.00) 4.00(0.00) 0.00(0.00) 1.46(0.05) -5.00(1.03) 4.54(0.01)
Table 6: Performance of Different Models in LLM as Independent System 1 and System 2 Experiments.
@@
ReAct Reflexion DPT-Agent DPT-Agent ReAct Reflexion DPT-Agent DPT-Agent
@@ I.1 Procedure
In the formal experiment (Figure 8), after each game, participants were asked to rank the agents based on their collaborative capabilities and personal preference. Upon completing all eight games, they filled out an additional questionnaire, where we collected their perceptions of task load and their intended level of task engagement.